### v1

@@ Abstract
Index Terms— Multi-channel, speech separation, sound source localization, mask estimation, speaker-independent
1 Introduction
@@
While popular concepts like permutation-invariant training (PIT) [6] or its multi-channel extension location-based training (LBT) [7] provide an efficient training framework, they are conceptually limited to a predetermined upper bound on simultaneous speech sources. This dependence arises from the fixed output size, which is deeply rooted in the architecture of the NN. While iterative separation approaches can circumvent this issue by extracting each speaker at a time in a spatially guided fashion [2, 4, 8, 9, 10, 11], the amount of subsequent NN executions in scenarios with many participants limits their applicability [12]. Neural sound source localization (SSL) systems [13] tackle the closely related problem of speaker-independent localization by assuming that all speakers are uniquely identifiable through their relative positioning towards the microphone array. By partitioning the recording environment into a fixed spatial grid, such approaches conduct either a binary or probabilistic speaker activity estimation for each individual region [14, 15]. Based on the success in SSL, this method has been adapted for speech separation. Instead of an activity indication, mask-weighted spatial binary coding (MW-SBC) encodes a time-frequency mask into each partition containing a speaker [16, 17, 18, 19]. However, in case of a very fine grid and sparse speech masks, there arises a significant label imbalance leading to an ill-conditioned optimization problem during training with regression-based loss functions [16, 18]. To avoid this issue, an alternative is to only utilize the output channels corresponding to the positions of the speakers. While this significantly improves the conditioning, it alleviates the need for the NN to conduct precise localization, thus a separate SSL system becomes necessary [18].
@@
2 Beamformer-guided Speech Separation
Let the speech mixture contain microphone recordings of stationary speech sources in a reverberant but noiseless environment. The signal propagation between -th source and all microphones shall be represented by the room impulse response (RIR) . In the short-time Fourier transform (STFT) domain at time frame and frequency bin , the acoustic scenario is resembled by
Due to its robustness and straightforward parameterization, we choose the minimum-variance distortionless response (MVDR) beamformer [20] for speech separation. Assuming knowledge about the microphone array geometry, it can be specified through the steering vector and covariance matrix . Dropping indices for visual clarity, the separated speech signals are given by
@@ 2 Beamformer-guided Speech Separation
3 Mask-Weighted Spatial Coding
3.1 Mask-Weighted Spatial Binary Coding
@@
with denoting the Kronecker delta [24], a binary indicator function which is non-zero only if equals . This can be seen as spectral extension of the binary DoA encoding [2, 25, 26, 27] which we will denote by spatial binary coding (SBC). In case a high spatial resolution is desired, the number of segments is much greater than the number of speakers . Consequently MW-SBC suffers from a significant label imbalance. Throughout the training of a NN, the resulting vast amount of zeros has a significant impact on the gradient during backpropagation. Especially regression-based loss functions, which are known to be robust against outliers, are prone to converge to poor, quasi-stationary solutions for the estimate of , such as [23]. Due to being a famous representative of regression-based loss functions, we will examine the mean squared error (MSE) more closely. Omitting the average about time and frequency bins, the MSE between and is defined by
@@
| as it weights all differentiations w.r.t. the NN parameters due to the chain rule. Its norm regarding the spatial dimension yields | ||||
with representing the spatial vectorization of . Examining the norm at and inserting (5) gives
@@
Assuming the remaining differential chain is bounded, the sub-multiplicative property of norms [28] leads to a vanishing of the whole gradient w.r.t. the NN parameters. In other words, with a very large amount of segments , is correct in almost all cases. The remaining non-zero partitions containing IRMs only have negligible influence, which leads to a plateau in training. To avoid this problem, [16, 18] solely utilize the masks of active speakers during loss computation. While this removes the majority of zeros, it also decouples precise localization from the training objectives.
3.2 Proposed Mask-Weighted Spatial Likelihood Coding
@@
with controlling the width of the Gaussians. The angular distance represents the wrapped mean absolute error (MAE) [32], thus taking the circularity of the DoAs into account. Similar to [15], we accumulate the individual curves with a maximum operation. However, in our case, due to weighting with the IRMs, the peaks of the underlying curves are no longer equally prominent. Especially for small angular distances and frequency bins dominated by a single speaker, adjacent masks can influence each other in our coding. While a chosen in correspondence to a minimum spatial gap between the speakers can alleviate this effect, it is a conceptual limitation MW-SBC does not suffer. To enable further analysis, we approximate the non-linear maximum operator in (11) with a sum
which is valid under the assumption of sharp and remote Gaussians. Based on this approximation, we conduct the same investigation for the MSE gradient norm at as for MW-SBC. Combining (8) with (11) at yields
| We expand this expression with the constant | ||||
with representing the angular space around the array in which the speakers can be located, e.g. in case of unconstrained speaker positioning . Considering the limiting case of a very high spatial resolution, thus the amount of segments approaches infinity, the partition size transitions to the differential . Together with the sum over the infinite amount of segments, the limit converges to the integral expression
Since the Gaussian function is strictly positive, the integral also evaluates to a positive non-zero value. Therefore, in contrast to MW-SBC in (10), the gradient’s norm of our proposed MW-SLC depends on the IRMs and is non-zero in speech presence. The main difference between both approaches lies in the continuous nature of the Gaussian encoding, which increases the spatial prominence with a rising amount of partitions , while it stays constant for the Kronecker delta function used in (5). To find a closed-form solution, we assume an unconstrained speaker positioning. By choosing the lower (equals upper) limit of integration centered between two DoAs, the circularity of the angular distance becomes negligible due to the assumption of remote and sharp Gaussians. In the same manner, the integration can be extended to the whole real number line, yielding |
The norm’s proportionality to underlines the strong connection between problem conditioning and width of the Gaussians. While an increased broadness improves training, it also puts less weight on accurate localization and possibly results in a reduced mask quality.
3.3 Joint DoA and Mask Extraction
@@
4 Experiments
4.1 Experimental Setup
Dataset For training and evaluation we utilize the publicly available multi-channel dataset MC-LibriMix [8]. Specifically, we employ the versions with a sampling rate of 16-kHz generated for two (MC-Libri2Mix) and three (MC-Libri3Mix) speakers. The spatialization was conducted with simulated RIRs for a linear 4-channel microphone array (details in [8]). Regarding the STFT, we use a square-root Hann window [33] of length 32 ms and 16 ms hop-size.
DoA and Mask Estimation The width of the Gaussian curves in both SLC and MW-SLC is set to 6 as in [8]. The same value is used for the neighborhood in (17) as proposed in [15]. The threshold is determined via an exhaustive search on the validation dataset of MC-Libri2Mix w.r.t. optimizing the score. Finally, we utilize hierachical agglomerative clustering (HAC) to first obtain the DoAs and then the IRMs for all identified speakers. In each step we use the average distance between separate clusters as linkage and set as merging-condition. The threshold for the IRM computation in (4) is set to -35 dB.
4.2 Network Architectures and Training Details
@@
5 Results
Table 1 displays the results regarding SSL and joint localization and mask estimation in terms of separation performance with an MVDR beamformer.

### v2

@@ Abstract
Index Terms:
Multi-channel, speech separation, sound source localization, mask estimation, speaker-independentI Introduction
@@
While popular concepts like permutation-invariant training (PIT) [6] or its multi-channel extension location-based training (LBT) [7] provide an efficient training framework, they are conceptually limited to a predetermined maximum amount of simultaneous speech sources. This dependency arises from the fixed output size, which is deeply rooted in the architecture of the NN. While iterative separation approaches can circumvent this issue by extracting each speaker at a time in a spatially guided fashion [2, 4, 8, 9, 10, 11], the amount of subsequent NN executions in scenarios with many participants limits their applicability [12]. Neural sound source localization (SSL) systems [13] tackle the closely related problem of speaker-independent localization by assuming that all speakers are uniquely identifiable through their relative positioning towards the microphone array. By partitioning the recording environment into a fixed spatial grid, such approaches conduct either a binary or probabilistic speaker activity estimation for each individual region [14, 15]. Based on the success in SSL, this method has been adapted for speech separation. Instead of an activity indication, mask-weighted spatial binary coding (MW-SBC) encodes a time-frequency mask into each partition containing a speaker [16, 17, 18, 19]. However, in the case of a very fine grid and sparse speech masks, there arises a significant label imbalance leading to an ill-conditioned optimization problem during training with regression-based loss functions [16, 18]. To avoid this issue, an alternative is to only utilize the output channels corresponding to the positions of the speakers. While this significantly improves the conditioning, it alleviates the need for the NN to conduct precise localization, thus a separate SSL system becomes necessary [18].
@@
II Beamformer-guided Speech Separation
Let the speech mixture contain microphone recordings of stationary speech sources in a reverberant but noiseless environment. The signal propagation between -th source and all microphones shall be represented by the room impulse response (RIR) . In the short-time Fourier transform (STFT) domain at time frame and frequency bin , the acoustic scenario is thus resembled by
Due to its robustness and straightforward parameterization, we choose the minimum-variance distortionless response (MVDR) beamformer [20] for speech separation. Assuming knowledge about the microphone array geometry, it can be specified through the steering vector and the covariance matrix . Dropping indices for visual clarity, the separated speech signals are given by
@@ II Beamformer-guided Speech Separation
III Mask-Weighted Spatial Coding
III-A Mask-Weighted Spatial Binary Coding
@@
with denoting the Kronecker delta [24], a binary indicator function which is non-zero only if equals . This can be seen as spectral extension of the binary DoA encoding [2, 25, 26, 27] which we will denote by spatial binary coding (SBC). In case a high spatial resolution is desired, the number of segments is much greater than the number of speakers , thus, the majority of regions are empty. Additionally, due to the sparsity of speech signals in the time-frequency domain, the bins of the remaining partitions containing speakers are also predominantly inactive, which is why MW-SBC suffers from a significant label imbalance. Throughout the training of a NN, the resulting vast amount of zeros has a significant impact on the gradient during backpropagation. Especially regression-based loss functions, which are known to be robust against outliers, are prone to converge to poor, quasi-stationary solutions for the estimate of , such as [23]. Due to being a famous representative of regression-based loss functions, we will examine the mean squared error (MSE) more closely. Omitting the average about time and frequency bins, the MSE between and is defined by
@@
| as it weights all differentiations w.r.t. the NN parameters according to the chain rule. Its norm regarding the spatial dimension yields | ||||
| with representing the spatial vectorization of . Examining the norm at and inserting (5) gives | ||||
@@
Assuming the remaining differential chain is bounded, the sub-multiplicative property of norms [28] leads to a vanishing of the whole gradient w.r.t. the NN parameters. In other words, with a very large amount of segments , is correct in almost all cases. The remaining non-zero partitions containing IRMs have only negligible influence, which leads to a plateau in training. To avoid this problem, [16, 18] solely utilize the masks of active speakers during loss computation. While this removes the majority of zeros, it also decouples precise localization from the training objectives.
III-B Proposed Mask-Weighted Spatial Likelihood Coding
@@
with the standard deviation of the Gaussians. The angular distance represents the wrapped mean absolute error (MAE) [32], thus taking the circularity of the DoAs into account. Similar to [15], we accumulate the individual curves with a maximum operation. However, in our case, due to weighting with the IRMs, the peaks of the underlying curves are no longer equally prominent. Especially for small angular distances and frequency bins dominated by a single speaker, adjacent masks can influence each other in our coding. While a chosen in correspondence to a minimum spatial gap between the speakers can alleviate this effect, it is a conceptual limitation MW-SBC does not suffer. To enable further analysis, we approximate the non-linear maximum operator in (12) with a sum
which is valid under the assumption of Gaussians with low standard deviations and large mean difference. Based on this approximation, we conduct the same investigation for the MSE gradient norm at as for MW-SBC. Combining (8) with (12) at gives
We rewrite this expression by introducing the constant
with representing the angular space around the array in which the speakers can be located, e.g. in the case of unconstrained speaker positioning . Considering the limiting case of a very high spatial resolution, thus the amount of segments approaches infinity, the partition size transitions to the differential . Together with the sum over the infinite amount of segments, the limit converges to the integral expression
Since the Gaussian function is strictly positive, the integral also evaluates to a positive non-zero value. Therefore, in contrast to MW-SBC in (11), the gradient’s norm of our proposed MW-SLC depends on the IRMs and is non-zero in speech presence. The main difference between both approaches lies in the continuous nature of the Gaussian encoding, which increases the spatial prominence with a rising amount of partitions , while it stays constant for the Kronecker delta function used in (5). To find a closed-form solution, we assume an unconstrained speaker positioning. By choosing the lower (equals upper) limit of integration centered between two DoAs, the circularity of the angular distance becomes negligible due to the assumption of Gaussians with low standard deviations and large mean difference. In the same manner, the integration can be extended to the whole real number line, yielding |
The norm’s proportionality to underlines the strong connection between problem conditioning and standard deviation of the Gaussians. While an increased improves training, it also puts less weight on accurate localization and possibly results in a reduced mask quality.
III-C Joint DoA and Mask Extraction
@@
IV Experiments
IV-A Experimental Setup
Dataset For training and evaluation we utilize the publicly available multi-channel dataset MC-LibriMix [8]. Specifically, we employ the versions with a sampling rate of 16-kHz generated for two (MC-Libri2Mix) and three (MC-Libri3Mix) speakers. The spatialization was conducted with simulated RIRs for a linear 4-channel microphone array. The simulations are constrained to reverberation times between 200 ms and 600 ms and a minimum angular distance between adjacent speakers of 15 ∘ (more details in [8]). Regarding the STFT, we use a square-root Hann window [33] of length 32 ms and 16 ms hop-size.
DoA and Mask Estimation The standard deviation of the Gaussian curves in both SLC and MW-SLC is set to 6 ∘ as in [8]. The same value is used for the neighborhood in (18) as proposed in [15]. The threshold is determined via an exhaustive search on the validation dataset of MC-Libri2Mix w.r.t. optimizing the score. Finally, we utilize hierachical agglomerative clustering (HAC) to first obtain the DoAs and then the IRMs for all identified speakers. In each step we use the average distance between separate clusters as linkage and set as merging-condition. The threshold for the IRM computation in (4) is set to -35 dB.
IV-B Network Architectures and Training Details
@@
V Results
Table I displays the results regarding SSL and joint localization and mask estimation in terms of separation performance with an MVDR beamformer. The loss computation is indicated by columns Encoding and Loss with SBC and SLC following [25] and [15] respectively.